OpenStudy (anonymous):

Which of the following is the equation for the line parallel to 2x + 5y = 15 and passes through the point (4, -4)?

OpenStudy (anonymous):

where r u ur options?

OpenStudy (anonymous):

5y=-2x+15 y=-2/5x+15

OpenStudy (anonymous):

can u tell me wt is the slope of dis equation?

OpenStudy (anonymous):

a. y - 2 = 4/3(x - 7) b. y + 4 = 4/5(x - 4) c. y + 7 = 4/3(x - 2) d. y + 4 = -2/5(x - 4)

OpenStudy (anonymous):

slope -2/5

OpenStudy (anonymous):

yes correct! when the lines are parallel m1=m2 dat means slope of rquird line will be same i.e. -2/5

OpenStudy (anonymous):

the requird eqn is, (y-y1)=m(x-x1)

OpenStudy (anonymous):

x1,y1 is the given point (4,-4) and slope is -2/5

OpenStudy (anonymous):

put all the values and u will get the eqn

OpenStudy (anonymous):

i need it in slope intercept form

OpenStudy (anonymous):

first u pu the values and match wid ur options

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

so wts the correct option?

OpenStudy (anonymous):

D

OpenStudy (anonymous):